OpenStudy (anonymous):

Does the formula: f(x) = 1/( x^2 – 2) define a function f: R → R A function f: Z → R

OpenStudy (anonymous):

it's a function as i think

OpenStudy (anonymous):

It is defined everywhere and there is no x in Z such that x^2=2

OpenStudy (anonymous):

The text is really weak in discussing the methods to solve. Should I just use arbitrary values to determine if the formula can be used for functions f: R to R and F; Z to R

OpenStudy (anonymous):

pretty much how i'd do it...^^^ same goes for \(\large f:Z\rightarrow R \) ... it's easy to see that there is no integer such that when you square that integer, you will get 2....

OpenStudy (anonymous):

I am slightly confused on performing the steps to verify the results. Why would I need to get a 2 for the solution?

OpenStudy (anonymous):

is that last line a separate question? or is it the same f as the one on the first?

OpenStudy (anonymous):

separate question

OpenStudy (anonymous):

I mean the f: Z to R applies to the question

OpenStudy (anonymous):

yeah, so in that second question, it's referring to the same function...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I probably need to show my work on this. So how would I show this

OpenStudy (anonymous):

the function f maps the integers to the reals... \(\large f:\mathbb{Z}\rightarrow\mathbb{R} \)